OpenStudy (anonymous):

local min/max of f(x)=ln(x^4+8) Did I do this correctly?

OpenStudy (anonymous):

\[f(x)=\ln(x^4+8)\] \[f'(x)=\frac{ 4x^3 }{ x^4+8 }\] Critical point would be 4x^3 = 0 so, x = 0 Next draw number line and check increasing/decreasing intervals:  because \[f'(-1)=\frac{ - }{ + }=negative\] \[f'(1)=\frac{ + }{ + }=positive\]

OpenStudy (anonymous):

oh, guess I need to find actual local min and max. So, I use my drawing to say, the local min is 0 and there is no local max for this equation? Is this correct?

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

now you just need to find the actual minimum value at x=0

OpenStudy (anonymous):

oh, that's what I'm forgetting. Thank you. Gonna figure that out.

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

so, plug x=0 into the original function.\[f(0)=\ln(0^4+8)\] \[f(0)=\ln(8)\] and I should've stated the increasing/decreasing intervals. Increasing on (0,infinity). Decreasing on (-infinity,0). How's that?

OpenStudy (asnaseer):

yes - you have it all perfectly correct. :) althought you should specifically mention that the minimum value = f(0) = ln(8)

OpenStudy (anonymous):

ok good. thank you!

OpenStudy (asnaseer):

yw :)